# On the Isbell Problem

### v1

@@ 3. $(iii)$
  4. (4)
    $p_{n}\Vdash({\tilde{p}_{n-1}}^{\frown}\tilde{\rho}_{n-1},\tilde{q}_{n-1}*\beta_{n})\in[\mathbf{P}^{(*\beta_{n})}*\dot{\mathbb{R}}:\mathbf{P}_{0}^{\beta_{n}}]\cap\mathcal{M}$.
  5. (5)
    $(p_{n},q_{n})\Vdash({\tilde{p}_{n-1}}^{\frown}\tilde{\rho}_{n-1},\tilde{\tau}_{n-1}^{\frown}\tilde{q}_{n-1})\in D_{m_{n-1}}$, where $m_{n-1}=\eta(n-1)$.
  6. (6)
@@
  7. (7)
    $p_{n}\Vdash\tilde{p}_{n-1}\Vdash\tilde{\rho}_{n-1}\leq\tilde{\varphi}_{2}(\gamma_{\Lambda(n-1)})$.
Let $\Gamma_{<\beta_{n}}$ be a $\mathbf{P}^{\beta_{n}}$-generic filter such that $(p_{n},q_{n})\in\Gamma_{<\beta_{n}}$ and let us work first in $V[\Gamma_{<\beta_{n}}^{0}]$. Then we have, $({\tilde{p}_{n-1}}^{\frown}\tilde{\rho}_{n-1},\tilde{q}_{n-1}*\beta_{n})\in[\mathbf{P}^{(*\beta_{n})}*\dot{\mathbb{R}}:\mathbf{P}_{0}^{\beta_{n}}]\cap\mathcal{M}$. Extend $\tilde{p}_{n-1}$ to a condition $\tilde{p}_{n}^{\prime}$ which decides the value of $\tilde{\varphi}_{2}(\gamma_{\Lambda(n)})$, that is, let $\tilde{r}_{n}\in dom(\dot{\mathbb{R}})$ be such that
$\tilde{p}_{n}^{\prime}\Vdash\tilde{\varphi}_{0}(\gamma_{\Lambda(n)})=\tilde{\gamma}_{2}(\gamma_{\Lambda(n)})=\tilde{r}_{n}$
Note that we have $\tilde{r}_{n}\in\mathcal{M}$. Let $m_{n}=\eta(n)$, so $n\in P_{m_{n}}$ and $n$ is the $(\Lambda(n))^{th}$-element of $P_{m_{n}}$. Consider the condition $({\tilde{p}_{n}^{\prime}}^{\frown}\tilde{r}_{n},\tilde{q}_{n-1}*\beta_{n})\in[\mathbf{P}^{(*\beta_{n})}*\dot{\mathbb{R}}:\mathbf{P}_{0}^{\beta_{n}}]\cap\mathcal{M}$. Since $\dot{E}_{m_{n}}^{n}$ is an open dense subset of $[\mathbf{P}^{(*\beta_{n})}*\dot{\mathbb{R}}:\mathbf{P}_{0}^{\beta_{n}}]\cap\mathcal{M}$, we can find $(p^{\frown}\dot{\rho},q)\in\dot{E}_{m_{n}}^{n}$ such that $(p^{\frown}\dot{\rho},q)\leq(({\tilde{p}_{n}^{\prime}}^{\frown}\tilde{r}_{n},\tilde{q}_{n-1}*\beta_{n})$. Let $({\tilde{p}_{n}}^{\frown}\tilde{\rho}_{n},\tilde{q}_{n})$ be a $\mathbf{P}_{0}^{\beta_{n}}$-name for $(p^{\frown}\rho,q)$. Now, going to $V[\Gamma_{<\beta_{n}}]$, since $({\tilde{p}_{n}}^{\frown}\tilde{\rho}_{n},\tilde{q}_{n})\in\dot{E}_{m_{n}}^{n}\subseteq(\overline{D}_{m_{n}})^{\divideontimes}$, we find $\tau$ such that:
  1. (1)
    $({\tilde{p}_{n}}^{\frown}\tilde{\rho}_{n},\tau^{\frown}\tilde{q}_{n})\in D_{m_{n}}$.
  2. (2)
@@
  1. (1)
    $\beta_{n}$ taking the place of $\beta_{0}$.
  2. (2)
@@
  3. (3)
    $[\mathbf{P}^{(*\beta_{n+1})}*\dot{\mathbb{R}}:\mathbf{P}_{0}^{\beta_{n+1}}]$ taking the place of $\dot{\mathbb{R}}$.
  4. (4)
@@
  5. (5)
    $(\tilde{p}_{n}\upharpoonright\beta_{n+1},\tilde{q}_{n}\upharpoonright\beta_{n+1})$ taking the place of $(\tilde{p},\tilde{q})$ and $({\tilde{p}_{n}}^{\frown}\tilde{\rho}_{n},\tilde{q}_{n}*\beta_{n+1})$ taking the place of $\tilde{r}$.
  6. (6)
    $\{\overline{D}_{k}^{\beta_{n+1}}:k\in\omega\}$ taking the place of $\{D_{k}:k\in\omega\}$, where $\overline{D}_{k}^{\\ beta_{1}}$ is defined as:
    $\overline{D}_{k}^{\beta_{n+1}}=\{(\Psi_{\beta_{n+1}}(p^{\frown}\dot{r},q):(p^{\frown}\dot{r},q)\in D_{k}\}$
Thus, we find $(p_{n+1},q_{n+1})\in\mathbf{P}^{\beta_{n+1}}$ which is $(\mathbf{P}_{0}^{\beta_{n+1}},\mathbf{P}^{\beta_{n+1}},\mathcal{M})$-generic for $\{\overline{D}_{k}^{\beta_{n+1}}:k\in\omega\}$ and $[\mathbf{P}^{(*\beta_{n+1})}*\dot{\mathbb{R}}:\mathbf{P}^{\beta_{n+1}}_{0}]$, and a family of $\mathbf{P}_{0}^{\beta_{n+1}}$-names $\{\dot{E}_{k}^{n+1}:k\in\omega\}$ such that:
  1. (1)
@@
  2. (2)
    $p_{n+1}\Vdash\dot{E}_{k}^{n+1}\subseteq[\mathbf{P}^{(*\beta_{n+1})}*\dot{\mathbb{R}}:\mathbf{P}_{0}^{\beta_{n+1}}]\cap\mathcal{M}$ is an open dense set.
  3. (3)
    $(p_{n+1},q_{n+1})\Vdash\dot{E}_{k}^{n+1}\subseteq(\overline{D}_{k}^{\beta_{n+1}})^{\divideontimes}$.
  4. (4)
    $(p_{n+1},q_{n+1})\Vdash(\tilde{p}_{n}\upharpoonright\beta_{n+1},\tilde{\tau}_{n}^{\frown}\tilde{q}_{n}\upharpoonright\beta_{n+1})\in\Gamma_{<\beta_{n+1}}\cap\mathcal{M}$.
  5. (5)
@@
  6. (6)
    $p_{n+1}\Vdash({\tilde{p}_{n}}^{\frown}\tilde{\rho}_{n},\tilde{q}_{n}*\beta_{n+1})\in[\mathbf{P}^{(*\beta_{n+1})}*\dot{\mathbb{R}}:\mathbf{P}_{0}^{\beta_{n+1}}]\cap\mathcal{M}$.
and additionally,
  1. (1)
    $p_{n+1}\Vdash\tilde{r}_{n}=\tilde{\varphi}_{2}(\gamma_{\Lambda(n)})$.
  2. (2)
@@
  3. (3)
    $p_{n+1}\Vdash({\tilde{p}_{n}}^{\frown}\tilde{\rho}_{n},\tilde{q}_{n})\in\dot{E}_{m_{n}}^{n}$, where $m_{n}=\eta(n)$.
  4. (4)
    $(p_{n+1},q_{n+1})\Vdash({\tilde{p}_{n}}^{\frown}\tilde{\rho}_{n},\tilde{\tau}_{n}^{\frown}\tilde{q}_{n})\in D_{m_{n}}$.
@@
    For each $n\in\omega$, $\tilde{F}_{n}=\{(\tilde{\rho}_{k},\overline{p}):k\in P_{n}\}$.
We claim that $(\overline{p},\overline{q})$ is $(\mathbf{P}_{0},\mathbf{P},\mathcal{M})$-generic for $\{D_{n}:n\in\omega\}$ and $\dot{\mathbb{R}}$.
@@
which implies $(p^{\prime}\upharpoonright\beta_{m},q^{\prime}\upharpoonright\beta_{m})\leq(p\upharpoonright\beta_{m},\tau^{\frown}q\upharpoonright\beta_{m})$, for each $m>k$. Finally note that $supp(p),supp(\tau^{\frown}q)\subseteq\bigcup_{l\in\omega}\beta_{l}$, since $(p,\tau^{\frown}q)\in\mathcal{M}$, so we can conclude that $(p^{\prime},q^{\prime})\leq(p,\tau^{\frown}q)$. Clearly this happens for any condition $(p^{\prime},q^{\prime})$ which determines $(\tilde{p}_{k},\tilde{q}_{k})$ and $\tilde{\tau}_{k}$, so we get $(\overline{p},\overline{q})\Vdash(\tilde{q}_{k},\tilde{\tau}_{k}^{\frown}\tilde{q}_{k})\in\dot{\Gamma}$.
Let us now see that each $\tilde{F}_{k}$ is forced by $(\overline{p},\overline{q})$ to be a dense subset of ˙$\dot{\mathbb{R}}\cap\mathcal{M}[\Gamma_{<\alpha}^{0}]$. Let $\Gamma$ be a $\mathbf{P}$-generic filter such that $(\overline{p},\overline{q})\in\Gamma$. Pick $r\in\dot{\mathbb{R}}_{\Gamma_{<\alpha}^{0}}\cap\mathcal{M}[\Gamma_{<\alpha}^{0}]$. Then there is $l\in\omega$ such that $r=\varphi_{0}(\gamma_{l})$. Note that there is $j\in P_{k}$ such that $\theta_{k}(j)=l$, so we have $\Lambda(j)=\theta_{k}(j)=l$. Thus, by construction we have,
@@
    $p_{j+1}\Vdash r=\tilde{r}_{j}=\tilde{\varphi}_{2}(\gamma_{\Lambda(j)})$.
@@
    $p_{j+1}\Vdash({\tilde{p}_{j}}^{\frown}\tilde{\rho}_{j},\tilde{q}_{j})\in\dot{E}_{k}^{j}$.
From $(\overline{p},\overline{q})\Vdash(\tilde{p}_{j},\tilde{\tau}_{j}^{\frown}\tilde{q}_{j})\in\Gamma$ we get $\overline{p}\Vdash\tilde{p}_{j}\in\Gamma_{<\alpha}^{0}$, so we conclude $(\tilde{\rho}_{j})_{\Gamma_{<\alpha}^{0}}\leq(\tilde{r}_{j})_{\Gamma_{<\alpha}^{0}}=r$. Finally note that $(\tilde{\rho}_{j})_{\Gamma_{<\alpha}^{0}}\in(\tilde{F}_{k})_{\Gamma_{<\alpha}^{0}}$. Thus, $(\tilde{F}_{k})_{\Gamma_{<\alpha}^{0}}$ is a dense subset of ˙$\dot{\mathbb{R}}_{\Gamma_{<\alpha}^{0}}\cap\mathcal{M}[\Gamma_{<\alpha}^{0}]$.
It remains to prove that $(\overline{p},\overline{q})\Vdash\tilde{F}_{k}\subseteq D_{k}^{\divideontimes}$. Pick $r\in(\tilde{F}_{k})_{\Gamma_{<\alpha}^{0}}$. By definition of $\tilde{F}_{k}$, there is $l\in P_{k}$ such that $r=(\tilde{\rho}_{l})_{\Gamma_{<\alpha}^{0}}$. Then we have $(\tilde{p}_{l},\tilde{\tau}_{l}^{\frown}\tilde{q}_{l})\in\Gamma$, and $({\tilde{p}_{l}}^{\frown}\tilde{\rho}_{l},\tilde{\tau}_{l}^{\frown}\tilde{q}_{l})\in D_{k}$, from which it follows that $r=(\tilde{\rho}_{l})_{\Gamma_{<\alpha}^{0}}\in D_{k}^{\divideontimes}$.
@@
If $\mathbf{P}$ is a $\varsigma$-proper iteration, then $(\mathbf{P}_{0},\mathbf{P})$ is $\varsigma$-proper.
@@
Throughout this section, $\mathbb{Q}_{\kappa}$ is the forcing from Definition 5.1. First let us see how the forcing $\mathbb{Q}_{\kappa}$ fits into the $\varsigma$-iterations framework.
@@
$\mathbb{Q}_{\kappa}$ is a $\varsigma$-iteration.
@@
It is well known that any countable support iteration of the Sacks forcing is strongly proper and has the Sacks property, so it is Cohen preserving as well. By Theorem 4.18, $\mathbb{P}_{\kappa}(\mathcal{F})$ has the Sacks property, so it is Cohen preserving; also, the proof of Theorem 4.18 can be easily adapted to prove that each $\mathbb{P}_{\kappa}(\mathcal{F})$ is strongly proper when $\mathcal{F}$ is a suitable filter. Therefore, any iteration of the form $\mathbb{S}_{\beta}*\dot{\mathbb{P}}_{\kappa}(\tilde{\mathcal{F}})$ is strongly proper and Cohen preserving, whenever $\tilde{\mathcal{F}}$ is forced by $\mathbb{S}_{\beta}$ to be a suitable filter.
Thus, we can define $\mathbb{Q}_{0}^{0}=\mathbb{S}_{\alpha_{0}}$ and $\mathbb{Q}_{0}^{1}=\mathbb{S}_{\alpha_{0}}*\dot{\mathbb{P}}_{\kappa}(\dot{\mathcal{U}}_{\alpha_{0}})$, and for $\beta+1$, ˙${\dot{\mathbb{Q}}}_{\beta+1}^{0}$and ˙${\dot{\mathbb{Q}}}_{\beta+1}^{1}$ are defined such that:
$\mathbb{S}_{\alpha_{\beta}}\Vdash\dot{\mathbb{Q}}_{\beta+1}^{0}=[\mathbb{S}_{\alpha_{\beta+1}}:\mathbb{S}_{\alpha_{\beta}}]$
and
$\mathbb{S}_{\alpha_{\beta}}\Vdash\dot{\mathbb{Q}}_{\beta+1}^{1}=[\mathbb{S}_{\alpha_{\beta+1}}*\dot{\mathbb{P}}_{\kappa}(\dot{\mathcal{U}}_{\alpha_{\beta+1}}):\mathbb{S}_{\alpha_{\beta}}]$
If $\beta$ is a limit ordinal we have two cases:
  1. Case 1.
    $\sup\{\alpha_{\gamma}:\gamma<\beta\}<\alpha_{\beta}$. Let $\delta_{\beta}=\sup\{\alpha_{\gamma}:\gamma<\beta\}$ and define,
    $\mathbb{S}_{\delta_{\beta}}\Vdash\dot{\mathbb{Q}}_{\beta}^{0}=[\mathbb{S}_{\alpha_{\beta}}:\mathbb{S}_{\delta_{\beta}}]$
    and
    $\mathbb{S}_{\delta_{\beta}}\Vdash\dot{\mathbb{Q}}_{\beta}^{1}=[\mathbb{S}_{\alpha_{\beta}}*\dot{\mathbb{P}}_{\kappa}(\dot{\mathcal{U}}_{\alpha_{\beta}}):\mathbb{S}_{\delta_{\beta}}]$
  2. Case 2.
    $\sup\{\alpha_{\gamma}:\gamma<\beta\}=\alpha_{\beta}$. Just define $\mathbb{S}_{\alpha_{\beta}}\Vdash\dot{\mathbb{Q}}_{\beta}^{0}=\{1\}$ and $\mathbb{S}_{\alpha_{\beta}}\Vdash\dot{\mathbb{Q}}_{\beta}^{1}=\dot{\mathbb{Q}}_{\beta}^{0}*\dot{\mathbb{P}}_{\kappa}(\dot{\mathcal{U}}_{\alpha_{\beta}})$.

### v2

@@ 5. $(5_{n})$
  6. $(6_{n})$
    $(p_{n},q_{n})\Vdash({\tilde{p}_{n-1}}^{\frown}\tilde{\rho}_{n-1},\tilde{\tau}_{n-1}^{\frown}\tilde{q}_{n-1})\in H_{n-1}$.
  7. $(7_{n})$
@@
  8. $(8_{n})$
    $p_{n}\Vdash\tilde{p}_{n-1}\Vdash\tilde{\rho}_{n-1}\leq\tilde{\varphi}_{2}(\gamma_{n-1})$.
Let $\Gamma_{<\beta_{n}}$ be a $\mathbf{P}^{\beta_{n}}$-generic filter such that $(p_{n},q_{n})\in\Gamma_{<\beta_{n}}$ and let us work first in $V[\Gamma_{<\beta_{n}}^{0}]$. Then we have, $({\tilde{p}_{n-1}}^{\frown}\tilde{\rho}_{n-1},\tilde{q}_{n-1}*\beta_{n})\in[\mathbf{P}^{(*\beta_{n})}*_{\mathbf{P}_{0}^{\alpha}}\dot{\mathbb{R}}:\mathbf{P}_{0}^{\beta_{n}}]\cap\mathcal{M}$. Extend $\tilde{p}_{n-1}$ to a condition $\tilde{p}_{n}^{\prime}\in[\mathbf{P}_{0}^{\alpha}:\mathbf{P}_{0}^{\beta_{n}}]\cap\mathcal{M}$ which decides the value of $\tilde{\varphi}_{2}(\gamma_{n})$, that is, let $\tilde{r}_{n}\in dom(\dot{\mathbb{R}})$ be such that
$\tilde{p}_{n}^{\prime}\Vdash\tilde{\varphi}_{0}(\gamma_{n})=\tilde{\varphi}_{2}(\gamma_{n})_{\Gamma_{<\beta_{n}}^{0}}=(\tilde{r}_{n})_{\Gamma_{<\beta_{n}}^{0}}$
Note that we can get $\tilde{r}_{n}\in\mathcal{M}$. Consider the condition $({\tilde{p}_{n}^{\prime}}^{\frown}\tilde{r}_{n},\tilde{q}_{n-1}*\beta_{n})\in[\mathbf{P}^{(*\beta_{n})}*_{\mathbf{P}_{0}^{\alpha}}\dot{\mathbb{R}}:\mathbf{P}_{0}^{\beta_{n}}]\cap\mathcal{M}$. Since $\dot{E}^{n}_{\Gamma_{<\beta_{n}}^{0}}$ is an open dense subset of $[\mathbf{P}^{(*\beta_{n})}*_{\mathbf{P}_{0}^{\alpha}}\dot{\mathbb{R}}:\mathbf{P}_{0}^{\beta_{n}}]\cap\mathcal{M}$, we can find $(p^{\frown}\dot{\rho},q)\in\dot{E}^{n}_{\Gamma_{<\beta_{n}}^{0}}$ such that $(p^{\frown}\dot{\rho},q)\leq({\tilde{p}_{n}^{\prime}}^{\frown}\tilde{r}_{n},\tilde{q}_{n-1}*\beta_{n})$. Let $({\tilde{p}_{n}}^{\frown}\tilde{\rho}_{n},\tilde{q}_{n})$ be a $\mathbf{P}_{0}^{\beta_{n}}$-name for $(p^{\frown}\rho,q)$. Now, going to $V[\Gamma_{<\beta_{n}}]$, since $({\tilde{p}_{n}}^{\frown}\tilde{\rho}_{n},\tilde{q}_{n})\in\dot{E}^{n}_{\Gamma_{<\beta_{n}}^{0}}\subseteq(\overline{H_{n}}^{\beta_{n}})^{\divideontimes}$, we find $\tau$ such that:
  1. $(A_{n})$
    $({\tilde{p}_{n}}^{\frown}\tilde{\rho}_{n},\tau^{\frown}\tilde{q}_{n})\in H_{n}$.
  2. $(B_{n})$
@@ 3. $(C_{n})$
  1. $i_{n})$
    $\beta_{n}$ taking the place of $\beta_{0}$ and $\beta_{n+1}$ taking the place of $\alpha$.
  2. $ii_{n})$
@@
  3. $iii_{n})$
    $[\mathbf{P}^{(*\beta_{n+1})}*_{\mathbf{P}_{0}^{\alpha}}\dot{\mathbb{R}}:\mathbf{P}_{0}^{\beta_{n+1}}]$ taking the place of $\dot{\mathbb{R}}$.
  4. $iv_{n})$
@@
  5. $v_{n})$
    $(\tilde{p}_{n}\upharpoonright\beta_{n+1},\tilde{q}_{n}\upharpoonright\beta_{n+1})$ taking the place of $\tilde{\eta}$.
  6. $vi_{n})$
    $\mathcal{H}_{\beta_{n+1}}(D,\mathbf{P},\dot{\mathbb{R}})$ taking the place of $\mathcal{H}_{\alpha}(D,\mathbf{P},\dot{\mathbb{R}})$.
Therefore, we find $(p_{n+1},q_{n+1})\in\mathbf{P}^{\beta_{n+1}}$ which is $(\mathbf{P}_{0}^{\beta_{n+1}},\mathbf{P}^{\beta_{n+1}},\mathcal{M})$-generic for $\mathcal{H}_{\beta_{n+1}}(D,\mathbf{P},\dot{\mathbb{R}})$ and $[\mathbf{P}^{(*\beta_{n+1})}*_{\mathbf{P}_{0}^{\alpha}}\dot{\mathbb{R}}:\mathbf{P}^{\beta_{n+1}}_{0}]$, so in particular there is a $\mathbf{P}_{0}^{\beta_{n+1}}$-name $\dot{E}^{n+1}$ such that:
  1. $\text{I}_{n}$)
    $p_{n+1}\Vdash\dot{E}^{n+1}\subseteq[\mathbf{P}^{(*\beta_{n+1})}*_{\mathbf{P}_{0}^{\alpha}}\dot{\mathbb{R}}:\mathbf{P}_{0}^{\beta_{n+1}}]\cap\mathcal{M}$ is an open dense set.
  2. $\text{II}_{n}$)
    $(p_{n+1},q_{n+1})\Vdash\dot{E}^{n+1}\subseteq(\overline{H_{n+1}}^{\beta_{n+1}})^{\divideontimes}$.
and we also have,
  1. $\text{III}_{n}$)
@@
  2. $\text{IV}_{n}$)
    $p_{n+1}\Vdash({\tilde{p}_{n}}^{\frown}\tilde{\rho}_{n},\tilde{q}_{n}*\beta_{n+1})\in[\mathbf{P}^{(*\beta_{n+1})}*_{\mathbf{P}_{0}^{\alpha}}\dot{\mathbb{R}}:\mathbf{P}_{0}^{\beta_{n+1}}]\cap\mathcal{M}$.
  3. $\text{V}_{n}$)
    There is a $\mathbf{P}^{\beta_{n}}$-name $\tilde{\tau}_{n}$ such that
    $(p_{n+1},q_{n+1})\Vdash(\tilde{p}_{n}\upharpoonright\beta_{n+1},\tilde{\tau}_{n}^{\frown}\tilde{q}_{n}\upharpoonright\beta_{n+1})\in\Gamma_{<\beta_{n+1}}\cap\mathcal{M}$
    and
    $(p_{n+1},q_{n+1})\Vdash({\tilde{p}_{n}}^{\frown}\tilde{\rho}_{n},\tilde{\tau}_{n}^{\frown}\tilde{q}_{n})\in H_{n}\subseteq D$
and additionally,
  1. $\text{VI}_{n}$)
@@
  2. $\text{VII}_{n}$)
    $p_{n+1}\Vdash\tilde{r}_{n}=\tilde{\varphi}_{2}(\gamma_{n})$.
  3. $\text{VIII}_{n}$)
@@
  4. $\text{IX}_{n}$)
    $p_{n+1}\Vdash({\tilde{p}_{n}}^{\frown}\tilde{\rho}_{n},\tilde{q}_{n})\in\dot{E}^{n}$.
@@ 4. $\text{IX}_{n}$)
    $\tilde{F}_{0}=\{(\tilde{\rho}_{k},\overline{p}):k\in\omega\}$.
  4. (4)
    $\tilde{F}_{1}^{l}=\{(\tilde{p}_{k}^{\frown}\tilde{\rho}_{k},\overline{p}):k\in\omega\land k\geq l\}$, for each $l\in\omega$.
We claim that $(\overline{p},\overline{q})$ is $(\mathbf{P}_{0},\mathbf{P},\mathcal{M})$-generic for $\mathcal{H}_{\alpha}(D,\mathbf{P},\dot{\mathbb{R}})$.
@@
which implies $(p^{\prime}\upharpoonright\beta_{m},q^{\prime}\upharpoonright\beta_{m})\leq(p\upharpoonright\beta_{m},\tau^{\frown}q\upharpoonright\beta_{m})$, for each $m>k$. Finally note that $supp(p),supp(\tau^{\frown}q)\subseteq\bigcup_{l\in\omega}\beta_{l}$, since $(p,\tau^{\frown}q)\in\mathcal{M}$, so we can conclude that $(p^{\prime},q^{\prime})\leq(p,\tau^{\frown}q)$. Clearly this happens for any condition $(p^{\prime},q^{\prime})$ which determines $(\tilde{p}_{k},\tilde{q}_{k})$ and $\tilde{\tau}_{k}$, so we get $(\overline{p},\overline{q})\Vdash(\tilde{p}_{k},\tilde{\tau}_{k}^{\frown}\tilde{q}_{k})\in\dot{\Gamma}$.
On the other hand, by definition of $H_{k}$, we have $(\tilde{p}_{k},\tilde{\tau}_{k}^{\frown}\tilde{q}_{k})\in D_{k}$ (recall the construction at the beginning of this section). Therefore, $(\overline{p},\overline{q})$ is $(\mathbf{P},\mathcal{M})$-generic, and this implies that $(\overline{p},\overline{q})$ forces ˙$\dot{\mathbb{R}}_{\Gamma_{<\alpha}^{0}}\cap\mathcal{M}[\Gamma_{<\alpha}^{0}]=\dot{\mathbb{R}}_{\Gamma_{<\alpha}^{0}}\cap\mathcal{M}[\Gamma]$. This will be assumed in the rest of the argument.
Let us now see that $\tilde{F}_{0}$ is forced by $(\overline{p},\overline{q})$ to be a dense subset of ˙$\dot{\mathbb{R}}\cap\mathcal{M}[\dot{\Gamma}_{<\alpha}^{0}]$. Let $\Gamma$ be a $\mathbf{P}$-generic filter such that $(\overline{p},\overline{q})\in\Gamma$. Pick $r\in\dot{\mathbb{R}}_{\Gamma_{<\alpha}^{0}}\cap\mathcal{M}[\Gamma_{<\alpha}^{0}]$. Then there is $j\in\omega$ such that $r=\varphi_{0}(\gamma_{j})$. Thus, by construction we have,
@@
    $p_{j+1}\Vdash r=\tilde{r}_{j}=\tilde{\varphi}_{2}(\gamma_{j})$.
@@
    $p_{j+1}\Vdash({\tilde{p}_{j}}^{\frown}\tilde{\rho}_{j},\tilde{q}_{j})\in\dot{E}^{j}$.
From $(\overline{p},\overline{q})\Vdash(\tilde{p}_{j},\tilde{\tau}_{j}^{\frown}\tilde{q}_{j})\in\Gamma$ we get $\overline{p}\Vdash\tilde{p}_{j}\in\Gamma_{<\alpha}^{0}$, so we conclude $(\tilde{\rho}_{j})_{\Gamma_{<\alpha}^{0}}\leq(\tilde{r}_{j})_{\Gamma_{<\alpha}^{0}}=r$. Finally note that $(\tilde{\rho}_{j})_{\Gamma_{<\alpha}^{0}}\in(\tilde{F}_{0})_{\Gamma_{<\alpha}^{0}}$. Thus, $(\tilde{F}_{0})_{\Gamma_{<\alpha}^{0}}$ is a dense subset of ˙$\dot{\mathbb{R}}_{\Gamma_{<\alpha}^{0}}\cap\mathcal{M}[\Gamma_{<\alpha}^{0}]$.
Let us prove that $(\overline{p},\overline{q})\Vdash\tilde{F}_{0}\subseteq(D)^{\divideontimes}$. Pick $r\in(\tilde{F}_{0})_{\Gamma_{<\alpha}^{0}}$. By definition of $\tilde{F}_{0}$, there is $l\in\omega$ such that $r=(\tilde{\rho}_{l})_{\Gamma_{<\alpha}^{0}}$. Then we have $(\tilde{p}_{l},\tilde{\tau}_{l}^{\frown}\tilde{q}_{l})\in\Gamma$, and $({\tilde{p}_{l}}^{\frown}\tilde{\rho}_{l},\tilde{\tau}_{l}^{\frown}\tilde{q}_{l})\in D$, from which it follows that $r=(\tilde{\rho}_{l})_{\Gamma_{<\alpha}^{0}}\in(D)^{\divideontimes}_{\Gamma}$. Now we verify clause (2) from the conclusions by proving that for each $l\in\omega$, $(\tilde{F}_{1}^{l})_{\Gamma_{<\alpha}^{0}}$ is predense in $[\mathbf{P}_{0}^{\alpha}*\dot{\mathbb{R}}:\mathbf{P}_{0}^{\alpha}]\cap\mathcal{M}[\Gamma]$ and $(\tilde{F}_{1}^{l})_{\Gamma_{<\alpha}^{0}}\subseteq(\overline{H_{l}}^{\alpha})^{\divideontimes}$. Assume $\Gamma$ is as before and pick an arbitrary condition $p^{\frown}\dot{r}\in[\mathbf{P}_{0}^{\alpha}*\dot{\mathbb{R}}:\mathbf{P}^{\alpha}_{0}]\cap\mathcal{M}[\Gamma]$, so we have $\dot{r}_{\Gamma_{<\alpha}^{0}}\in\dot{\mathbb{R}}_{\Gamma_{<\alpha}^{0}}\cap\mathcal{M}[\Gamma_{<\alpha}^{0}]$ and there is $n\geq l$ such that $(\tilde{\rho}_{n})_{\Gamma_{<\alpha}^{0}}\leq\dot{r}_{\Gamma_{<\alpha}^{0}}$, and actually, $(\tilde{p}_{n}^{\frown}\tilde{\rho}_{n},\tilde{\tau}_{n}^{\frown}\tilde{q}_{n})\in H_{l}$ (since $H_{n}\subseteq H_{l}$), which implies that $\tilde{p}_{n}^{\frown}\tilde{\rho}_{n}\in(\overline{H_{l}}^{\alpha})^{\divideontimes}$. Since $\tilde{p}_{n},p\in\Gamma_{<\alpha}^{0}$, we have that $\tilde{p}_{n}^{\frown}\tilde{\rho}_{n}$ and $p^{\frown}\dot{r}$ are compatible.
All it remains is to prove clause $(5)$: this follows from the facts $(\overline{p},\overline{q})\Vdash(\tilde{p}_{j},\tilde{\tau}_{j}^{\frown}\tilde{q}_{j})\in\dot{\Gamma}$, $(p_{1},q_{1})\Vdash(\tilde{p}_{0},\tilde{\tau}_{0}^{\frown}\tilde{q}_{0})\leq(\tilde{\eta}_{0},\tilde{\nu}^{\frown}\tilde{\eta}_{1})$ and $(\overline{p},\overline{q})\leq(p_{1},q_{1})$. ∎
@@
If $\mathbf{P}$ is a $\varsigma$-iteration of Cohen proper forcings, then $(\mathbf{P}_{0},\mathbf{P})$ is Cohen proper.
@@
Throughout this section, $\mathbb{Q}_{\kappa}$ is the forcing from Definition 5.1. First let us see how the forcing $\mathbb{Q}_{\kappa}$ fits into the Cohen proper iterations framework.
@@
$\mathbb{Q}_{\kappa}$ is a Cohen proper iteration.
@@
We follow the same notation from section 5. It is not hard to see that any pair of the form $(\mathbb{S},\mathbb{S}\times\mathbb{P}_{\kappa}(\mathcal{F}))$ is Cohen proper, whenever $\mathcal{F}$ is a suitable filter.
Thus, we can define $\mathbb{Q}_{0}^{0}=\mathbb{S}$ and $\mathbb{Q}_{0}^{1}=\mathbb{S}*1$, where $1$ is the trivial forcing. Now, for $\gamma\notin D_{0}$, we make $\mathbb{S}_{\gamma}\Vdash\dot{\mathbb{Q}}_{0}^{\gamma}=\dot{\mathbb{S}}$ and $\mathbb{S}_{\gamma}\Vdash\dot{\mathbb{Q}}_{1}^{\gamma}=\dot{\mathbb{S}}*1$. If $\gamma\in D_{0}$, let $\beta\in\omega_{2}$ be such that $\gamma=\alpha_{\beta}$. Then make $\mathbb{S}_{\alpha_{\beta}}\Vdash\dot{\mathbb{Q}}_{0}^{\alpha_{\beta}}=\dot{\mathbb{S}}$ and $\mathbb{S}_{\alpha_{\beta}}\Vdash\dot{\mathbb{Q}}_{1}^{\alpha_{\beta}}=\dot{\mathbb{S}}\times\dot{\mathbb{P}}_{\kappa}(\dot{\mathcal{U}}_{\alpha_{\beta}})^{V[G_{\alpha_{\beta}}]}$. It is easy to see that this iteration is isomorphic to $\mathbb{Q}_{\kappa}$. ∎